### v1

# The Social Life of Industrial Arms: How Arousal and Attention Shape Human-Robot Interaction

Abstract
This study explores how human perceptions of a non-anthropomorphic robotic manipulator are shaped by two key dimensions of behaviour: arousal, defined as the robot’s movement energy and expressiveness, and attention, defined as the robot’s capacity to selectively orient toward and engage with a user. We introduce a novel control architecture that integrates a gaze-like attention engine with an arousal-modulated motion system to generate socially meaningful behaviours. In a user study, we find that robots exhibiting high attention—actively directing their focus toward users—are perceived as warmer and more competent, intentional, and lifelike. In contrast, high arousal—characterized by fast, expansive, and energetic motions—increases perceptions of discomfort and disturbance. Importantly, a combination of focused attention and moderate arousal yields the highest ratings of trust and sociability, while excessive arousal diminishes social engagement. These findings offer design insights for endowing non-humanoid robots with expressive, intuitive behaviours that support more natural human-robot interaction.
Figure 1: Photos showing varying robot poses as a result of varying the arousal level of the proposed architecture controlling the emotional and social behaviours of a Universal Robots UR5e manipulator.
I Introduction
As robots become increasingly integrated into human environments—from homes and classrooms to healthcare and industrial settings—designing them to interact naturally and intuitively with people is becoming a central challenge in human-robot interaction (HRI). While much of robotics has focused on functional performance, there is growing recognition that social and emotional capabilities are equally important for fostering trust, collaboration, and long-term engagement. This work explores a fundamental question: Can robots be designed with character-like qualities that support their integration into human-centred environments? We are particularly interested in how robots can use behavioural cues—such as movement, posture, and attention—to project emotional and social presence, even in the absence of anthropomorphic features. By endowing robots with the ability to exhibit emotionally expressive behaviour, we aim to deepen human-machine interaction and improve the quality of user experience.
As emotional machines become more prominent, it is essential to consider not only the capability of robots to communicate emotions but also how humans perceive these expressions. For instance, the findings from [1] suggest that adaptive emotional robots can foster rapport and immediacy in educational settings, thus leading to improved learning outcomes. Furthermore, according to [2], establishing emotional bonds and providing consistent support are key aspects of long-term human-computer relationships, which can be applied to addressing social isolation and similar challenges.
In this study, we focus on non-anthropomorphic robots (NARs) —specifically, robotic manipulators—and examine how they can express emotional character through two key behavioural dimensions: arousal, defined as the robot’s level of energy or movement intensity, and attention, defined as its capacity to selectively orient toward and engage with human users. Our proposed system architecture combines a simplified arousal model, grounded in Russell’s Circumplex Model of Affect [3], with an interactive focus mechanism we call the Attention Engine. To evaluate this architecture, we conduct a user study examining how variations in arousal and attention affect human perceptions of the robot’s warmth, competence, animacy, and social intentionality. The results offer insights into how expressive motion and attentional behaviours can enhance the social presence of NARs.
II Related Work
Traditionally, the ability to convey emotions has been reserved for more anthropomorphic robots– machines designed with human-like features that naturally lend themselves to the expression of human emotions [4]. Anthropomorphic robots have been widely studied in social human-robot interaction (HRI), particularly for their ability to use both verbal and non-verbal communication modes like kinesics, proxemics, haptics, and voice modulation [5, 6, 7]. Studies on such robots emphasize that nonverbal behaviours, including gestures and postures, can shift human cognitive framing, elicit emotional responses, and improve task performance [8]. Furthermore, eye gaze plays a critical role in human-robot communication, irrespective of the robot’s form. It has been shown to direct user attention, regulate conversational flow, and convey engagement or intent. Eye gaze has been used in multi-modal interaction to combine attention-directing cues with other communication modes, improving the robot’s ability to integrate into human environments [9, 10]. However, anthropomorphic robots also face social hurdles, such as the uncanny valley, a phenomenon in which a certain level of human likeness elicits feelings of eeriness and discomfort [11].
Instead, our work explores the use of NARs to behave socially and emotionally. NARs, such as robotic manipulators, lack human-like attributes, which makes the task of emotional expression different and more challenging. NARs often appear mechanical and devoid of personal content, lacking the emotional depth observed in their anthropomorphic counterparts [12]. However, prior work on social and emotional content designed for NARs discusses how postural adjustments in robots can shift emotional framing and trigger specific human responses [13, 14, 15]. Recent work has shown that expressive movement can significantly improve robot readability and engagement; studies incorporating animation principles, such as anticipation and reaction, demonstrate that motion design influences how users perceive a robot’s competence and intelligence [16]. Additionally, a movement-centric approach suggests that robot motion should be designed as a primary communication channel rather than an auxiliary feature [17]. Beyond animation-driven readability, the interplay between expressive and functional movement has been studied in NARs. The ELEGNT framework proposes that robots should integrate both function-driven and expression-driven movements to enhance user engagement and communication, even in robots designed primarily for task execution [18]. This work aligns with research showing that expressive motion helps users anticipate robot actions and infer internal states, making interactions more intuitive.
Our work aims to extend and explore how manipulator-based NARs can generate emotional contexts and social interactions with users. It appears that prior studies have not extensively investigated how robotic manipulators can emulate social behaviours typically associated with anthropomorphic designs, presenting a significant gap that this research addresses. In this research, we draw on emotional states framed in terms of Russell’s Circumplex Model of Affect [3], examining how a robot can convey “arousal” based on adjustments on a manipulator’s posture and motion intensity.
Arousal, which refers to the intensity of an emotional state, has been extensively linked to the perception of a robot’s energy, urgency, and emotional engagement. Higher arousal is often associated with increased activity levels, potentially signalling excitement or urgency, whereas lower arousal may convey calmness, disinterest, or boredom [19]. Given that emotional arousal plays a fundamental role in human emotional processing and response, studying its impact on the perception of robotic behaviours provides valuable insights into designing robots that elicit desired emotional reactions from users.
Attention, on the other hand, is critical in establishing engagement and social presence. A robot’s ability to direct attention towards a user conveys social awareness and intentionality, influencing how participants perceive the robot’s competence and sociability [20]. Exploring the effects of attention modulation, therefore, provides critical insights into how robotic systems can enhance user engagement and build rapport with their human counterparts.
Integrating movement-based attention and arousal mechanisms extends the scope of emotional communication in robots, highlighting the potential of universal, form-independent designs.
II-A Contributions
In this work, we present a behaviour generation architecture that enables emotional and social expression in NARs through the two chosen key behavioural dimensions: arousal and attention. Our contributions are as follows:
  • We introduce a novel control architecture that integrates an arousal-modulated motion model with a real-time attention engine that simulates gaze and focus behaviours in a robotic manipulator.
  • We conduct a factorial user study examining how variations in arousal and attention affect human perceptions of the robot’s warmth, competence, animacy, sociability, intentionality, disturbance, and discomfort.
  • We provide empirical evidence that while attention increases positive social perceptions, excessive arousal can reduce sociability and trust—highlighting the importance of balance in affective behaviour design for NARs.
These contributions are framed around the following research questions:
  1. 1.
    How can NARs convey emotional character through movement and interactivity without relying on anthropomorphic features?
  2. 2.
    What are the key factors in robotic movement and behaviours that can be optimized to evoke specific emotional responses from humans and enhance their emotional engagement?
  3. 3.
    How do varying levels of arousal and attention influence human perceptions of a robot’s social presence and intentionality?
III System Design
Figure 2: System architecture diagram for our proposed interactive and expressive robot behaviour system.
This section outlines the development of a behaviour generation system that enables NARs to interactively attend to people and environmental stimuli. As shown in Fig. 2, the system consists of two main components: an attention engine, which determines which stimuli the robot should attend to (Section III-A), and a gaze-tracking algorithm (Section III-B). We implemented this system on a Universal Robots UR5e manipulator [21], illustrated in Fig. 1, and evaluated it through a user study examining human perceptions of the robot (Section IV).
III-A Attention Engine
Our attention engine builds on the approach introduced by Pan et al. [22], originally developed for a humanoid animatronic bust. It governs the robot’s decision-making by computing an attention score for each detected individual, determining saliency and guiding the robot’s gaze to enable dynamic, responsive interactions.
Each attention score is a weighted combination of positional cues, movement dynamics, and a habituation factor that prevents prolonged fixation on a single target—reflecting how humans distribute attention across multiple people and stimuli. The score $\Phi$ is defined as:
$\Phi=w_{p}P+w_{v}V+\Theta(t)$ (1)
where $w_{p}$ and $w_{v}$ represent the weight for the position and velocity components, respectively. The weights are chosen based on their relative importance and influence on attention. $\Theta(t)$ is a habituation factor, described in further detail below. The position score $P$ combines proximity and hand position factors:
$P=w_{proximity}e^{\lambda d}+w_{hand}(h_{left}+h_{right})$ (2)
where $d$ is the Euclidean distance to the individual, $\lambda$ is a decay constant, and $h_{\text{left}}$, $h_{\text{right}}$ are binary indicators of whether either hand is raised above shoulder level. This formulation prioritizes nearby individuals and interprets raised hands as social gestures, such as waving. $\lambda$ is used to modulate the influence of distance on the attention score, ensuring that closer individuals are prioritized, similarly to natural human tendency.
The velocity score $V$ captures torso and hand motion, normalized by maximum observed movement:
$V=\frac{v_{torso}}{v_{max\_torso}}+\frac{v_{right}}{v_{max\_right}}+\frac{v_{left}}{v_{max\_left}}$ (3)
Movement is a key indicator of engagement, and normalization ensures responsiveness across varying activity levels.
To avoid sustained attention on a single user, we apply a habituation factor $\Theta(t)$ that dynamically adjusts based on gaze history:
$\Theta(t)=\Theta(t-1)+(\gamma m_{hab}+(1-\gamma)m_{rest})\Delta t$ (4)
Here, $\gamma=1$ for the currently attended individual and 00 for others; $m_{\text{hab}}$ and $m_{\text{rest}}$ control the decay and recovery rates, and $\Delta t$ is the time step. $\Theta(t)$ is bounded between 0 and 1 to maintain naturalistic attention shifts and prevent fixation.
III-B Robotic Gaze Algorithm
The robotic gaze algorithm transforms the 3D Cartesian position of the highest-scoring attention target into a joint-space posture for the manipulator, modulated by the system’s internal arousal level. While implemented on a UR5e arm in this study, the algorithm is designed to generalize across articulated platforms by prioritizing expressive, segmental motion patterns inspired by animation principles [23, 22]. Specifically, it favours movement in distal, lower-inertia joints over proximal ones—analogous to how gaze in animals and humans is typically guided by the head and neck rather than the whole body.
Arousal is treated as a continuous parameter ranging from 1 (lowest) to 10 (highest), influencing both the robot’s posture and movement dynamics. Lower arousal values produce subdued, grounded configurations with minimal motion, while higher values induce elevated, extended postures with increased responsiveness and motion amplitude.
The base joint (e.g., shoulder yaw or first revolute joint) adjusts laterally in proportion to the target’s horizontal displacement. This angle is constrained within safe operational limits and provides coarse orientation toward the target. Mid-chain joints (e.g., shoulder pitch and elbow) encode the manipulator’s forward or backward lean. These are influenced both by arousal and the target’s depth relative to the base frame—leaning forward for distant targets and backward for nearby ones under high arousal conditions. End-effector–adjacent joints provide finer tracking by calculating angular offsets that orient the robot’s terminal link toward the target in 3D space, accounting for current posture and reachability constraints.
To enhance perceived naturalness, the system overlays a subtle, sinusoidal oscillation on the joint angles to simulate breathing. The amplitude and frequency of this micro-motion scale with arousal—higher arousal leads to faster, shallower oscillations reflecting physiological excitement.
When a salient secondary target is detected—such as a raised hand—the gaze priority temporarily shifts to that target, creating the impression of a glance before returning to the primary attention focus. In the absence of human targets, the algorithm transitions into an idle mode where it generates virtual gaze targets within the workspace. These are tracked using the same joint-space computation and breathing modulation, preserving consistent motion characteristics across active and idle states.
III-C Virtual Object Manager
While the attention engine and gaze algorithm enable responsive tracking of human participants, early pilot studies revealed that continuous fixation on a user can quickly lead to interactions that feel repetitive or overly mechanical. This observation is consistent with prior work noting that overly persistent gaze in social robots can reduce perceived naturalness and engagement [24].
To address this, we introduce a Virtual Object Manager that injects naturalistic variability into the robot’s gaze behaviour. Inspired by human patterns of attentional distractibility [25], the system periodically generates virtual gaze targets within the robot’s workspace. These targets are designed to transiently divert the robot’s attention, simulating behaviours such as curiosity or environmental scanning.
Virtual targets are instantiated probabilistically, with the likelihood of generation increasing alongside the robot’s arousal level. Each target is assigned a randomized location in 3D space and a short-lived attention score that can briefly surpass those of human participants, prompting a shift in gaze. After a randomized lifespan, the target expires and is removed from consideration. This mechanism enables momentary deviations from human tracking while preserving consistency with the robot’s broader behavioural architecture.
By integrating virtual objects into the attention pipeline, the robot exhibits subtle gaze shifts that break monotony and promote more dynamic and lifelike interactions—even during periods of low human activity or idle behaviour.
IV User Study
To evaluate the effectiveness of the proposed behaviour generation system, we conducted a user study investigating how variations in arousal and attention influence human perceptions of a non-anthropomorphic robot. Building on the system described in Section III, this study examines the social and emotional impact of robot behaviour through open-ended human-robot interactions and was approved by the Queen’s University General Research Ethics Board (GELEC-139-22, File No. 6036728).
We specifically explore two behaviour dimensions: arousal, operationalized as the robot’s movement energy (cf. Russell’s Circumplex Model of Affect [3]), and attention, defined as the robot’s capacity to direct gaze and responsiveness toward users. These factors were selected due to their central role in shaping user perceptions of social agents, as discussed in Section II. A $2\times 2$ factorial design was employed, crossing two levels of arousal (low and high) with two levels of attention (low and high), resulting in four experimental conditions.
IV-A Conditions
  • Low Attention: The robot remains in an idle state and does not track or respond to human users. It intermittently shifts its gaze toward virtual stimuli generated by the virtual object manager.
  • High Attention: The robot actively tracks and responds to the participant with gaze and posture adjustments, following the mechanisms described in Sections III-A and III-B.
  • Low Arousal: Arousal is set to its minimum value (1), resulting in slow, conservative motion, a hunched posture, and limited range of movement to create a subdued and deliberate behavioural profile.
  • High Arousal: Arousal is set to its maximum value (10), producing faster, more dynamic movement with an upright posture and extended reach, conveying a heightened energy level.
IV-B Protocol
The study was conducted in a controlled lab environment. At the beginning of the experiment, participants were asked to fill out a pre-study survey based on the General Attitude Towards Robots Survey (GAToRS) [26], which assessed participants’ pre-existing perceptions and attitudes toward robots, providing a baseline for understanding the audience’s general outlook towards robots. The survey was administered as a series of 7-point Likert-type questions with response options ranging from “Strongly Disagree” (1) to “Strongly Agree” (7).
During the study, participants were invited to engage with the robot in an interaction session without a specific time limit, allowing them to end the interaction whenever they chose and to freely explore the robot’s interactive and dynamic behaviours. Although they were instructed not to have any part of the body cross into the robot’s workspace (denoted by a line on the floor), they could attempt to interact with the robot in any way they wished.
Each participant experienced four interaction sessions corresponding to the four conditions of the experiment. To control for potential carryover effects between conditions, a balanced Latin square was employed to determine the presentation order of the experimental conditions across participants. This ensured an even distribution of potential sequence effects.
After completing each experimental condition, participants were asked to fill out a combined version of the Robotic Social Attributes Scale (RoSAS) [27] and the Human-Robot Interaction Evaluation Scale (HRIES) [28]. This post-condition survey provided a comprehensive assessment of participants’ social perceptions of the robot, measuring the following constructs: warmth, competency, discomfort, sociability, agency, animacy, and disturbance.
All interactions were video and audio recorded with the consent of participants, enabling subsequent qualitative analysis. These recordings were intended to capture nuanced behavioural and verbal cues during the interactions and to offer potential interaction metrics, such as the length of the interactions under different experimental conditions. We were unable to capture video data from two participants; one did not provide consent and another due to technical difficulties.
V Results
A total of 36 participants were recruited through on- and off-campus advertising. Participation was voluntary and uncompensated. The sample included 18 males and 18 females (none identified as non-binary, transgender, or other gender identities), with ages ranging from 18 to 25 years ($M=21.7$, $SD=1.58$).
V-A Baseline Attitudes
Pre-study attitudes, as measured by the GAToRS survey [26], revealed generally positive perceptions of robots. On the personal level, participants reported moderate comfort and enjoyment around robots ($P^{+}=4.44$, $SD=1.52$) and low levels of unease ($P^{-}=2.93$, $SD=1.67$). On the societal level, participants expressed optimism about the impact of robots ($S^{+}=5.92$, $SD=1.05$), though concerns about potential risks remained present ($S^{-}=5.11$, $SD=1.60$), with more variability in responses.
V-B Main Effects and Interaction Analysis
A $2\times 2$ within-subjects repeated-measures ANOVA (RM-ANOVA) was conducted on the seven dependent variables measured by the combined RoSAS–HRIES instruments, as well as on interaction durations. Assumptions of sphericity and normality were not violated. Partial eta squared ($\eta^{2}$) is reported as a measure of effect size. Significant results are summarized in Table I, and interaction effects are visualized in Fig. 3.
Attention had a significant effect on all dependent measures except discomfort and disturbance. Arousal significantly affected discomfort, sociability, and disturbance. Notably, both discomfort and sociability also exhibited significant interaction effects between arousal and attention.
V-C Post Hoc Comparisons
Bonferroni-corrected pairwise comparisons revealed the following ratings out of seven:
[…]
Interaction Effects
[…]
TABLE I: Significant Results for Arousal and Attention Effects
Source Measure F Partial $\eta^{2}$ Power
Arousal Discomfort 10.703** .234 .889
Sociability 4.753* .120 .564
Disturbance 7.578** .178 .763
Attention Warmth 16.971*** .327 .980
Competence 175.349*** .834 1.000
Sociability 7.217** .171 .743
Animacy 18.051*** .340 .985
Intentionality 65.315*** .651 1.000
Arousal × Attention Discomfort 4.637* .117 .553
Sociability 5.376* .133 .616
[…]
Figure 3: Interaction effect of arousal and attention on discomfort and sociability ratings.
V-D Interaction Time
Interaction duration (in seconds) was also analyzed. Average times by condition were:
  • Low Arousal, Low Attention: $M=75.70$, $SD=33.51$
  • Low Arousal, High Attention: $M=68.20$, $SD=32.23$
  • High Arousal, Low Attention: $M=62.97$, $SD=26.50$
  • High Arousal, High Attention: $M=61.44$, $SD=38.34$
[…]
VI Discussion
The findings of this study highlight the importance of both arousal and attention in shaping users’ perceptions of non-anthropomorphic robotic manipulators. Significant arousal/attention interaction effects were detected for the RoSAS measure of discomfort and the HRIES measure of sociability, indicating that for low arousal conditions, the factor of attention does not appear to have a significant effect on reports of discomfort; whereas for high arousal conditions, high attention conditions significantly increase discomfort over the low attention conditions. In terms of sociability, it appears that for low arousal conditions, a robot demonstrating high attention was perceived to be more sociable than low attention, whereas, for high arousal conditions, sociability of the high attention conditions dropped, whereas reports for low attention remained relatively unchanged. Our results also indicate that the robot’s display of attention has a statistically significant effect on the RoSAS measures of warmth and competence, as well as the HRIES measures of intentionality and animacy. Variations in arousal predominantly affected feelings of disturbance.
High attention conditions yielded higher ratings of warmth, competence, and intentionality. When the robot was actively engaged with participants, its behaviours were interpreted as deliberate and controlling, suggesting a higher degree of agency. This conclusion is supported by participant feedback such as, “The robot seemed to follow my movement and then periodically look away, creating the sense that it was almost shy.” Others described it as “Cautiously curious. …Almost like a friendly dog.” These observations emphasize that when the robot “pays attention”, it is seen as more socially engaging and capable.
Increased arousal levels were associated with higher ratings of disturbance and discomfort. High arousal conditions, characterized by faster, more erratic movements, elicited feelings of unease among participants. One participant noted: “It made me feel uneasy since the tracking felt very intentional. It definitely felt like it was watching me,” while another remarked: “Clearly more aggressive. The fast-tracking made it seem hostile.” These responses indicate that while arousal can signal engagement and energy, excessive arousal may override positive social cues and trigger discomfort.
The interaction between arousal and attention further nuances our understanding of social perceptions. Under low arousal conditions, increasing attention appeared to enhance sociability and reduce discomfort. Participants in the low arousal, high attention condition described the robot as “It felt like it was tracking my head/eyes and trying to make a social connection,” suggesting that a calm demeanour combined with focused attention fosters trust and friendliness. However, when high arousal was paired with high attention, the resulting heightened energy increased discomfort and reduced sociability. One participant noted,“ This interaction made it feel as if it was an animal watching me in the wild,” while another mentioned, “It immediately felt very uncanny and creepy.” This finding suggests that the positive effects of high attention can be compromised when arousal is too intense, revealing the need for a balanced calibration.
The analysis of interaction times revealed that participants engaged for longer durations in low arousal conditions compared to high arousal conditions, suggesting that calmer robot behaviour encourages more sustained interactions. This finding aligns with the observed increase in discomfort and disturbance ratings in higher arousal conditions, potentially shortening participant engagement.
While our study provides valuable insights, several limitations must be acknowledged. First, the participant sample consisted mainly of university students within a narrow age range (18-25 years old), limiting the generalizability of the findings to a broader population. Second, our study primarily focused on short-term, freeform interactions ranging from three seconds to a few of minutes per condition. Interpretability of the findings may be limited as participants’ goals, motivations, and expectations were unconstrained. Additionally, It remains unclear how these perceptions evolve over prolonged engagement with the robot. Lastly, although the proposed system was described as platform-agnostic, all study data were collected using a UR5e manipulator; thus, participants may have attributed behaviours differently based on the form factor (industrial, mechanical arm), colouring perceptions.
Our findings suggest that the proposed attention and arousal-based system architecture effectively enables NARs to convey social and emotional states through movement and interaction. The attention engine’s ability to dynamically track and respond to human presence, combined with the arousal-modulated movement patterns, creates a foundation for socially engaging robotic behaviours. However, the results also indicate several areas for system refinement. The attention engine could be enhanced by incorporating more real-world environmental stimuli such as ambient noise levels, lighting conditions, and spatial context to better direct gaze and modulate attention, moving beyond the current virtual object-based approach. Additionally, the arousal system could be dynamically modulated based on environmental cues and interaction history, allowing the robot to transition through different emotional states that reflect its “experience” and context. These potential improvements, combined with the current findings, provide a clear direction for future iterations of the system architecture.
VII Conclusion
This work presents a behaviour generation system for NARs that leverages arousal and attention to support emotionally expressive and socially engaging interactions. Through a controlled user study, we examined how variations in these two dimensions influence human impressions of a robotic manipulator. Our findings show that high attention significantly enhances perceptions of warmth, competence, intentionality, and animacy, while high arousal—expressed through fast and expansive motion—can increase discomfort and disturbance. These results underscore the importance of balancing energy and engagement: attentional cues are effective in eliciting positive social responses, but excessive arousal may diminish the robot’s social acceptability.
The study demonstrates that even robots without human-like features can convey character-like behaviour when designed with appropriate expressive strategies. These insights offer concrete guidance for the design of socially intuitive robotic systems, particularly in non-humanoid forms. While our findings are limited by factors such as the short-term nature of the interactions and scope, this work lays the groundwork for future studies on long-term engagement, adaptive behaviours, and deployment in diverse real-world contexts.